# Leveraging Atom Loss Errors in Fault Tolerant Quantum Algorithms

### v2

@@ 5 Native loss detection from logical teleportation
A key observation is that a wide range of known logical subroutines naturally employ teleportation, inherently keeping qubit lifecycles short. Fig. 6(b) illustrates the average number of SE rounds per physical qubit, from initialization to measurement, across various essential subroutines. The results show that most algorithms have relatively brief lifecycles. Detailed descriptions of these algorithms are provided in the Supplementary Materials [50] Section S3. For this analysis, we conservatively assume one SE round per logical operation. However, as shown in Fig. 5(c), this number can be further reduced. Additionally, Fig. 3(d) and Supplementary Figure S11 [50] reveal that, at an experimentally-motivated loss fraction of 0.5, for fewer than 10 SE rounds, performance remains nearly unchanged without any active loss correction added to the SE rounds, relying solely on SSR and delayed-erasure decoding. These findings suggest that while memory benchmarks provide useful performance metrics, transversal logical algorithms are inherently well-equipped to manage loss, with minimal to no loss detecting SE. Leveraging SSR and delayed-erasure decoding can significantly enhance the performance of the logical algorithm without incurring additional experimental complexity.
@@
Figure 7: Deep logical algorithms with teleported gates. (a) We study a deep circuit consisting of teleported $X$ and $Z$ logical gates, similar in structure to small-angle synthesis algorithms [35, 36]. (b) Circuit-level simulation results for 11 layers of random teleported $Z$ and $X$ logical gates, with all logical qubits initialized in the presence of noise (physical error rate $p=1\%$ and $d=7$). The delayed-erasure decoder (pink) substantially outperforms an MLE decoder that does not account for loss (black) and improves with increasing loss fraction. It also matches the performance of a decoder with perfect loss information (gray, dashed), and for loss fractions less than 1, approaches that of the erasure channel (khaki, dashed). The erasure channel performance is achieved using Free SE with period 0.25, corresponding to loss detection and qubit replacement after every gate, thus providing a lower bound.
@@
As a proof of concept, we consider the teleportation-based logical circuit shown in Fig. 7(a), which mirrors the structure of the small-angle synthesis algorithm [35, 36]. This algorithm constructs small-angle rotations through sequences of $H$ and $T$ gates, with the latter often realized via teleportation. Consequently, loss detection and correction are integrated into the process of executing the logical algorithm. The logical error rates for the circuit in Fig. 7(a) are plotted in Fig. 7(b) as a function of the loss fraction using various decoders: a delayed-erasure decoder leveraging SSR, a regular MLE decoder, and an MLE decoder with perfect loss location information. The delayed-erasure decoder demonstrates significant performance improvements, with algorithmic logical error rates rapidly decreasing as the loss fraction increases, achieving the performance of the decoder with perfect loss information. Furthermore, for loss fractions less than 1, it achieves the lower bound set by the erasure channel—corresponding to loss detection and qubit replacement after every gate. Notably, this circuit is similar to the teleportation-based SE method investigated here in the memory setting, both utilizing SSR and replacing atoms frequently through teleportation without additional overhead. These results highlight the advantages of exploiting native loss detection within teleportation gadgets in logical circuits while also utilizing loss information to improve decoding. Finally, while our simulations focus on Clifford gate teleportation, they are expected to extend to non-Clifford gates, as the underlying loss detection principles remain consistent.
@@ 6 Conclusion
These considerations provide a framework for managing atom loss in neutral atom QEC, and explore the role of loss detection and bias in improving logical algorithm performance. Central to our approach is the use of the delayed-erasure decoder, which leverages loss information to approximate an MLE decoding solution, substantially improving the logical error rate.
Using our decoder, we examine how the algorithmic structure influences the optimal strategy for detecting and correcting loss. For high-depth circuits involving a large number of gate layers prior to logical measurement, the performance depends on both loss decoding and frequent loss detection and replacement. Using a surface code logical memory, we explore different SE methods under various loss fractions. Our findings indicate that, with appropriate decoding strategies, all methods enhance performance with increasing loss fraction. Notably, teleportation-based SE is a promising candidate for neutral atom quantum computing, as it leverages SSR to achieve comparably high thresholds for high loss fractions, albeit with an additional space overhead. Determining the optimal SE method for a given system will depend on experimental validation and the specific noise characteristics of the hardware.
By applying these results to multi-qubit deep logical algorithms, several key insights emerge. Atom loss errors, when managed using loss detection operations and SSR detection, are fully compatible with correlated decoding, allowing for an optimal number of approximately four logical operations per SE round using the delayed-erasure MLE decoder. This framework ensures effective error correction even in the presence of high loss rates.
@@
An intriguing avenue for future exploration involves optimizing loss detection in specific algorithmic subroutines, tailoring the loss detection frequency and SE method to align with the algorithmic gadget and system characteristics. Additionally, the delayed-erasure decoder can be naturally applied to the decoding of transversal non-Clifford circuits, as it is inherently compatible with and effective for correcting Clifford errors propagating through the circuit. Finally, our decoding techniques are broadly applicable to a wide range of QEC codes and logical algorithms, making them well-suited for further exploration of loss decoding in high-rate low-density parity-check (qLDPC) codes  [67, 68, 69, 70, 71, 72, 73, 74, 75].
Note added. Since the completion of our work, we became aware of several related studies examining the effect of qubit loss on a single logical memory: teleportation-SE [22], loss detection units [23], and SWAP-SE [76]. While addressing similar challenges, these works differ from ours in focus and methodology. Notably, improved thresholds have been recently reported in updated versions of Refs.[22, 76], based on the use of decoding strategies similar to the delayed erasure decoder described in this work. These contributions also utilize error models slightly different from those presented here. For completeness, our updated Appendix A includes results obtained using the error models utilized in Refs.  [22, 76].
@@ Acknowledgments
Appendix A SE Methods Lifecycles and Space-time Overheads
Category Conventional SWAP period 1 Free period 2 Teleportation-based Free period 1 Free period 1 + exact loss info Free period 0.25 + exact loss info
Space overhead $\sim 2d^{2}-1$ $\sim 2d^{2}-1$ $\sim 2d^{2}-1$ $\sim 3d^{2}-1$ $\sim 2d^{2}-1$ $\sim 2d^{2}-1$ $\sim 2d^{2}-1$
Time overhead $\sim 4d$ $\sim 4d$ $\sim 4d$ $\sim 6d$ $\sim 4d-1$ $\sim 4d-1$ $\sim 4d-1$
Threshold for $L=1$ [%] (Channel 2) NA $1.94\pm 0.01$ $2.75\pm 0.07$ $3.5\pm 0.05$ $3.76\pm 0.01$ $5.12\pm 0.01$ $6.5\pm 0.01$
Table  1 summarizes the average lifecycles, error thresholds, and space-time overheads for each SE method.
@@ Appendix B Delayed-Erasure Decoder
In this section, we will describe our approach for decoding in the presence of loss. We first describe the exact MLE solution. Then, we describe the approximated MLE decoding which we use in the main text.
@@ B.1 Exact MLE Decoding in the Presence of Loss
In this section, we provide a formulation of the most-likely-error (MLE) decoding problem in the presence of loss errors. Loss acts in a much more complex fashion compared to normal Pauli errors, and therefore require more care in terms of their formulation. The results also depend on the details of the error model and measurement model, introducing further complications. We will attempt to describe the full formulation in this section, which will then guide the choice of heuristic methods, primarily based on reweighting edges in the decoding graph, that will approximate the MLE problem.
We formulate the problem by generalizing the discussion in Ref. [12]. We can associate a binary variable $E_{i}$ for each Pauli error, and similarly $L_{i}$ for each loss event. We assume that the loss and Pauli errors are applied independently (i.e. the channel does not simultaneously apply two types of errors, even though that could be the case when doing a CZ gate, where there is a correlation between loss on one qubit and a Pauli error on the other). When running the circuit, we obtain information about errors and loss through qubit measurements. This information comes in two flavors: First, there are flags $F_{j}$, which indicate a qubit was lost when we tried to measure it. Second, there are detectors $D_{j}$, which are products of stabilizer measurement results that should be deterministic in the absence of errors. A non-trivial aspect of loss decoding is that the detectors must be determined based on the flags, since some qubit measurement results are labeled as “invalid” instead of 0 or 1.
@@
The error probability $P(\vec{E},\vec{L})$ can be readily written down based on the log likelihood ratios of individual events. One potential subtlety that one may need to worry about is that for a given qubit at a given lifecycle, only one loss event can occur. Thus, if we have a sequence of gates with loss probability $p_{1}$, $p_{2}$, …, then the loss event probabilities may need to be written as $p_{1}$, $(1-p_{1})p_{2}$, $(1-p_{1})(1-p_{2})p_{3}$,…, although the deviations may be higher-order in practice.
The more challenging factor to evaluate is $P(\vec{F},\vec{D}|\vec{E},\vec{L})$. First, we need to specify how the new detectors are chosen given a heralded loss pattern. In Ref. [42], they compare the stabilizer value against the next available syndrome measurement result at the same site. This seems like a reasonable strategy, since this bit of information is really not accessible until the next time we measure this stabilizer, and is not revealed by any other stabilizers. Note that strictly speaking, if we solve the full MLE problem, the precise choice of detector shouldn’t matter due to stabilizer equivalences, so long as it forms a complete basis. However, it will matter if we are using heuristic decoders. After forming the detectors, we need to evaluate whether a given loss and error pattern is consistent with the observed loss flags and detectors. The loss flag check is straightforward, and can be imposed as a constraint that exactly one of the loss locations corresponding to a given qubit are activated. The detector is more subtle, because now some of the way they are triggered becomes probabilistic. A brute force way to check would be as follows: For a given loss pattern, the circuit is now known. We can forward propagate all error events through the new circuit, and obtain measurement results, which will then tell us the probability distribution and allow us to calculate the conditional probability. Unfortunately, this seems rather costly to perform and may not be as easily phrased as a simple optimization problem with a linear optimization target, so we need to think a bit more about better ways to evaluate this probability.
Finally, we comment on the benefits and downsides of heuristic edge re-weighting techniques. For example, in Ref. [42], given a leakage event, they assign an $X$ and $Z$ error on the $i$th location with probability $i/2n$ and propagate this error forward to modify the edge weights. Note that the detailed error event is related to their leakage model, where all qubits that interact with a leaked qubit experience large depolarizing noise. This is good to leading order, since it captures the fact that qubits that interact with the measure qubit later are more likely to have experienced leakage. However, it seems to still miss one potentially important aspect: If a qubit is leaked at some point, then it is leaked for the rest, so the error events are actually correlated. This correlation can in principle be leveraged to more accurately deduce the error that occurred.
Given these considerations, the main goal is thus to find good heuristic methods that approximate MLE. In particular, some better intuition about the types of error events that are incorrectly classified for a given weighting could help inform modifications that make it closer to MLE. One could also either try different heuristics and see what gives the best logical error rate (focusing perhaps on “hard” instances to make the search more efficient), or implement MLE so that it also becomes clear how big the optimality gap to MLE is for a given heuristic (this also allows one to more clearly focus on concrete instances of heuristic decoding failures). With the possible exception of neural network decoders that automatically handle both aspects (see Google’s paper, where there are actually also simulations with heralded leakage [5]), it seems overall that closing the gap to MLE may make a bigger difference than the MLE-MLD gap.

### v3

@@ 5 Native loss detection from logical teleportation
A key observation is that a wide range of known logical subroutines naturally employ teleportation, inherently keeping qubit lifecycles short. Fig. 6(b) illustrates the average number of SE rounds per physical qubit, from initialization to measurement, across various essential subroutines. The results show that most algorithms have relatively brief lifecycles. Detailed descriptions of these algorithms are provided in Appendix I. For this analysis, we conservatively assume one SE round per logical operation. However, as shown in Fig. 5(c), this number can be further reduced. Additionally, Fig. 3(d) and Fig. 22 reveal that, at an experimentally-motivated loss fraction of 0.5, for fewer than 10 SE rounds, performance remains nearly unchanged without any active loss correction added to the SE rounds, relying solely on SSR and delayed-erasure decoding. These findings suggest that while memory benchmarks provide useful performance metrics, transversal logical algorithms are inherently well-equipped to manage loss, with minimal to no loss detecting SE. Leveraging SSR and delayed-erasure decoding can significantly enhance the performance of the logical algorithm without incurring additional experimental complexity.
@@
Figure 7: Deep logical algorithms with teleported gates. (a) We study a deep circuit consisting of teleported $X$ and $Z$ logical gates, similar in structure to small-angle synthesis algorithms [35, 36]. (b) Circuit-level simulation results for 11 layers of random teleported $Z$ and $X$ logical gates, with all logical qubits initialized in the presence of noise (physical error rate $p=1\%$ and $d=7$). The delayed-erasure decoder (pink) substantially outperforms an MLE decoder that does not account for loss (black) and improves with increasing loss fraction. It also matches the performance of a decoder with perfect loss information (gray, dashed), and for loss fractions less than 1, approaches that of the erasure channel (khaki, dashed). The erasure channel performance is achieved using direct conversion SE with period 0.25, corresponding to loss detection and qubit replacement after every gate, thus providing a lower bound.
@@
As a proof of concept, we consider the teleportation-based logical circuit shown in Fig. 7(a), which mirrors the structure of the small-angle synthesis algorithm [35, 36]. This algorithm constructs small-angle rotations through sequences of $H$ and $T$ gates, with the latter often realized via teleportation. Consequently, loss detection and correction are integrated into the process of executing the logical algorithm. The logical error rates for the circuit in Fig. 7(a) are plotted in Fig. 7(b) as a function of the loss fraction using various decoders: a delayed-erasure decoder leveraging SSR, a regular MLE decoder, and an MLE decoder with perfect loss time-location information. The delayed-erasure decoder demonstrates significant performance improvements, with algorithmic logical error rates rapidly decreasing as the loss fraction increases, achieving the performance of the decoder with perfect loss information. Furthermore, for loss fractions less than 1, it achieves the lower bound set by the erasure channel—corresponding to loss detection and qubit replacement after every gate. Notably, this circuit is similar to the teleportation-based SE method investigated here in the memory setting, both utilizing SSR and replacing atoms frequently through teleportation without additional overhead. These results highlight the advantages of exploiting native loss detection within teleportation gadgets in logical circuits while also utilizing loss information to improve decoding. Finally, while our simulations focus on Clifford gate teleportation, they are expected to extend to non-Clifford gates, as the underlying loss detection principles remain consistent.
@@ 6 Conclusion
These results provide a general framework for managing qubit loss or leakage errors across diverse quantum computing platforms, primarily neutral atoms but also including superconducting qubits and trapped ions, and highlight the impact of loss detection and bias on logical algorithm performance. Our analysis across a wide range of loss fractions and detection strategies reveals how loss detection and decoding can be leveraged in realistic regimes, and was already employed in the recent experimental studies of architectural mechanisms for universal fault-tolerant quantum computing using neutral atom arrays [10].
Central to our approach is the use of the delayed-erasure decoder, which leverages loss information to approximate an MLE decoding solution, substantially improving the logical error rate. Using our decoder, we examine how the algorithmic structure influences the optimal strategy for detecting and correcting loss. For high-depth circuits involving a large number of gate layers prior to logical measurement, the performance depends on both loss decoding and frequent loss detection and replacement. Using a surface code logical memory, we explore different SE methods under various loss fractions. Our findings indicate that, with appropriate decoding strategies, all methods enhance performance with increasing loss fraction. Notably, teleportation-based SE is a promising candidate for neutral atom quantum computing, as it leverages SSR to achieve comparably high thresholds for high loss fractions, albeit with an additional space overhead. Determining the optimal SE method for a given system will depend on experimental validation and the specific noise characteristics of the hardware.
By applying these results to multi-qubit deep logical algorithms, several key insights emerge. Loss errors, when managed using loss detection operations and SSR detection, are fully compatible with correlated decoding, allowing for an optimal number of approximately four logical operations per SE round using the delayed-erasure MLE decoder. This framework ensures effective error correction even in the presence of high loss rates.
@@
An intriguing avenue for future exploration involves optimizing loss detection in specific algorithmic subroutines, tailoring the loss detection frequency and SE method to align with the algorithmic gadget and system characteristics. Additionally, the delayed-erasure decoder can be naturally applied to the decoding of transversal non-Clifford circuits, as it is inherently compatible with and effective for correcting Clifford errors propagating through the circuit.
Finally, although our simulations focus on the surface code, the core techniques presented in this work are broadly applicable to a wide range of QEC codes and logical algorithms. In particular, our delayed-erasure decoder is agnostic to the underlying code and can be applied to any stabilizer code with a suitable decoding backend. Moreover, several loss-handling strategies explored here, such as erasure conversion and teleportation-based syndrome extraction, are compatible with a wide class of codes, including high-rate qLDPC codes [64, 65, 66, 67, 68, 69, 70, 71, 72].
Note added. Since the completion of our work, we became aware of several related studies examining the effect of qubit loss on a single logical memory: teleportation-SE [23], loss detection units [24], and SWAP-SE [73]. While addressing similar challenges, these works differ from ours in focus and methodology. Notably, improved thresholds have been recently reported in updated versions of Refs.[23, 73], based on the use of decoding strategies similar to the delayed erasure decoder described in this work. These contributions also utilize error models slightly different from those presented here. For completeness, our Appendices includes results obtained using the error models utilized in Refs.  [23, 73].
@@ Acknowledgments
Appendix A Syndrome of Loss Errors
This section explains how heralded loss alters the stabilizer measurement structure and modifies the decoding hypergraph in QEC codes. Each loss mechanism induces a modified circuit, which alters the set of stabilizer detectors and introduces new correlated error configurations. The resulting decoding hypergraph consists of a set of detectors (vertices) and hyperedges, where each hyperedge represents a possible error event, connects a subset of detectors it flips, and is assigned a corresponding probability. This hypergraph serves as the input to the decoder. Given a heralded loss at measurement (e.g., using SSR or erasure conversion), we reconstruct the lifecycle of the affected qubit (defined as the sequence of operations from initialization to measurement) and enumerate its potential loss locations. Each potential loss point defines a distinct “loss circuit,” which produces a different set of hyperedges in the decoding hypergraph.
As discussed in the main text, loss requires two decoder-level adjustments: (1) modifying detectors to remove dependence on the lost measurement (via superchecks), and (2) accounting for error edges arising from truncated circuits after loss. Below, we elaborate on each adjustment and provide a specific example for clarity.
First, in each loss circuit, the measurement of the lost qubit is set to a random result. This results in assigning a probability of $0.5$ to the relevant edge, which corresponds to a weight of 0 in the decoding hypergraph. This process generates a supercheck operator, defined as the product of neighboring checks, effectively eliminating the lost qubit’s contribution [51]. Since the loss is detected in all loss circuits for the potential loss locations, the supercheck operator appears in all decoding hypergraphs of the qubit’s lifecycle. Consequently, the supercheck is incorporated into the final decoding hypergraph of the lifecycle.
Second, when an atomic qubit is lost, subsequent gates involving this qubit act trivially. Therefore, each loss circuit omits certain gates due to the lost qubit. As a result, the decoding hypergraph for the circuit includes new edges that correspond to these missing gates, capturing the errors associated with the specific loss under consideration. These edges appear in the final decoding hypergraph of the lifecycle, with probabilities reflecting the likelihood of the associated events.
Figure 8: Example of detector activations generated by different loss events in a $d=3$ surface code over four time steps (t=1t=1–4). Panels (a) and (b) show two potential loss locations for the same data qubit $q$, in round $t=2$ and $t=3$, respectively, each after the first gate. These correspond to different loss circuits within the same lifecycle and are combined into a single decoding hypergraph for that lifecycle (labeled “lifecycle syndrome”). Panel (c) shows a loss of a measure qubit $m^{\prime}$ in round $t=2$, which belongs to a separate lifecycle and contributes a separate decoding hypergraph. Detectors $D_{0}$$D_{4}$ are highlighted in color. Each row ($E_{i}$) in the decoding hypergraph tables corresponds to one error mechanism (hyperedge), i.e., a correlated set of detector activations triggered by a single loss event. Superchecks that are activated in all loss circuits of a given lifecycle also appear in the final decoding hypergraph, such as the shared activation of $D_{3}$ in both (a) and (b).
Figure 8 illustrates a specific example using a $d=3$ surface code over four time steps: an initialization round ($t=1$), two syndrome extraction (SE) rounds ($t=2,3$), and a final transversal measurement ($t=4$). In this example, we present the syndrome patterns caused by losses at different locations and times. The upper panels show the circuits with canceled gates and heralding, while the lower panels display the decoding hypergraph, listing which detectors are activated and their correlations. This example uses a conventional SE scheme with measure qubits read out every round and data qubits only at the end, but our decoding approach applies to any SE method, as discussed in the main text.
Detectors $D_{0}$-$D_{4}$ correspond to parity checks between stabilizer outcomes across rounds: $D_{0}$ and $D_{1}$ compare the same stabilizer across $t=1\!\to\!2$ and $t=2\!\to\!3$ respectively, $D_{2}$ monitors the neighboring stabilizer of $q$ across $t=2\!\to\!3$, $D_{3}$ is the final plaquette stabilizer containing $q$ given by the final transversal measurement and compared with previous round, and $D_{4}$ compares measure qubit $m$’ across $t=1\!\to\!2$.
Each row in the decoding hypergraph (Figure 8) corresponds to a distinct detector activation pattern produced by propagating a specific loss through the truncated circuit. For example, when data qubit $q$ is lost at $t=2$, all subsequent CZ gates on $q$ are canceled and the final readout is randomized. The missing CZs make stabilizers involving $q$ inconsistent across rounds, producing time-like detector flips such as $\{D_{0},D_{1}\}$ or $\{D_{1}\}$. In some branches, the loss heralding itself introduces a plaquette mismatch ($D_{3}$), or the neighboring stabilizer of $q$ registers a time-like inconsistency ($D_{2}$). Thus the multiple rows for a single loss represent different deterministic ways a single heralded loss can violate stabilizer checks, depending on timing and Pauli by-product branches. Now we explain each row explicitly.
(a) Loss of data qubit $q$ at $t=2$. All subsequent gates on $q$ are canceled, and the final measurement is probabilistic. The missing CZs disrupt stabilizers involving $q$. The decoding hypergraph contains five rows: $\{D_{0},D_{1}\}$ (mismatches in both $t=1\!\to\!2$ and $t=2\!\to\!3$), $\{D_{1}\}$ (only $t=2\!\to\!3$), $\{D_{1},D_{3}\}$ (round $t=2\!\to\!3$ plus the plaquette check containing $q$), $\{D_{2}\}$ (neighbor stabilizer at $t=2\!\to\!3$), and $\{D_{3}\}$ (plaquette $t=3\!\to\!4$ inconsistency).
(b) Loss of data qubit $q$ at $t=3$. Since stabilizers remain intact through $t=2$, only the final checks are affected: $\{D_{1}\}$ (round $t=2\!\to\!3$), $\{D_{1},D_{3}\}$ (round $t=2\!\to\!3$ plus plaquette $t=3\!\to\!4$), and $\{D_{3}\}$ (plaquette $t=3\!\to\!4$ inconsistency).
Since both (a) and (b) correspond to different loss locations within the same data qubit lifecycle, any detectors that are activated in all loss circuits become part of the supercheck structure for that lifecycle. In this case, detector $D_{3}$, which corresponds to the final plaquette stabilizer involving $q$, is activated in both loss circuits and thus appears in both decoding hypergraphs. As a result, $D_{3}$ is retained in the final decoding hypergraph for the lifecycle and serves as a supercheck.
(c) Loss of measure qubit $m$’ at $t=2$. The round-2 measurement of $m$’ is missing, invalidating detectors $D_{2}$ and $D_{4}$. Their product forms a valid supercheck, $D_{2}D_{4}=M_{m’,t=1}M_{m’,t=3}$, correlating rounds 1 and 3 directly. Additionally, the canceled CZ introduces an error on the paired data qubit, activating $D_{0}$.
These examples show how a single loss modifies the circuit and produces specific correlated detector patterns.
@@ Appendix B Delayed-Erasure Decoder
This section presents our decoding strategy for handling heralded loss in fault-tolerant circuits. Loss acts in a much more complex fashion compared to normal Pauli errors, and therefore requires more care in terms of their formulation. Here, we show how to formulate the full MLE problem, and then construct a scalable decoder that approximates this solution.
@@ B.1 Exact MLE Decoding in the Presence of Loss
This section formulates the MLE problem in the presence of heralded qubit loss. While exact MLE decoding is computationally challenging, it serves as a theoretical benchmark that guides the development of efficient approximate methods.
We generalize the framework of Ref. [13] by incorporating both Pauli and loss errors. Let $\vec{E}=(E_{1},E_{2},\ldots)$ represent binary variables for Pauli error events, and $\vec{L}=(L_{1},L_{2},\ldots)$ represent binary variables for loss events. We assume Pauli and loss errors occur independently across circuit locations, although correlated models can be incorporated.
Measurement outcomes provide two kinds of information: (1) Flags $\vec{F}$, indicating when qubit loss is detected; and (2) Detectors $\vec{D}$, constructed as products of stabilizer measurement outcomes that are expected to be +1+1 in the absence of error. Because loss renders some measurements invalid, the set of active detectors must be dynamically redefined based on the observed flags.
@@
The error probability $P(\vec{E},\vec{L})$ can be calculated based on the log likelihood ratios of individual events. However, since only one loss event can occur per qubit lifecycle, the effective probability of loss at location $j$ becomes $p_{j}\prod_{k<j}(1-p_{k})$. This correction is relevant when the loss probabilities vary significantly across locations.
The more challenging factor to evaluate is $P(\vec{F},\vec{D}|\vec{E},\vec{L})$. First, given a loss pattern, detectors must be updated by replacing invalidated measurements with combinations of adjacent ones (e.g., superchecks). While the exact choice is irrelevant in full MLE decoding due to stabilizer equivalence, it can affect the performance of heuristic decoders. After forming the detectors, we need to evaluate whether a given loss and error pattern is consistent with the observed loss flags and detectors. The flag check is straightforward: at most one loss event should be active per qubit lifecycle. Evaluating the detector consistency is more subtle. In the presence of loss, certain detector activations become probabilistic, depending on both the location and timing of the loss event. Given a specific loss pattern, the circuit becomes fixed, and the resulting conditional detector distribution can in principle be computed by propagating the error configuration through the modified circuit. However, this approach is computationally expensive and difficult to express as a simple optimization problem, motivating approximate methods.
One such approach is edge reweighting, where known correlations between loss and downstream Pauli errors are captured by modifying decoder edge weights. For instance, Ref. [42] models leaked qubits as inducing depolarizing noise on all subsequent interactions, assigning higher error weights to gates further in time. While this captures leading-order effects, it neglects correlations introduced by the fact that leaked qubits persist across time steps, resulting in temporally correlated error patterns. Accounting for this structure can improve heuristic approximations to MLE decoding.
Ultimately, the goal is to identify practical heuristic methods that approximate MLE decoding while capturing the effects of loss. This can be pursued by analyzing which errors are commonly misclassified by a given weighting scheme or by benchmarking against exact MLE decoding on small instances. Neural network decoders offer a complementary strategy by directly learning mappings from syndrome and flag data to logical corrections, implicitly capturing both error and loss correlations [5]. However, for near-term systems where clarity and integration with standard decoders are important, improving heuristic approximations to MLE may offer greater practical benefits.